Hi!
You invest $15,000 in an account with an interest rate of 8%. If interest accrues continuously how long will it take for your investment to double?
thanks for the study help!

Answer by Alan3354(69443)

You invest $15,000 in an account with an interest rate of 8%. If interest accrues continuously how long will it take for your investment to double?
--------------
For compounding continuously (I think that's what you mean by accruing):
FV = Amt*e^(rt)
e^0.08t = 2
0.08t = ln(2)
t = ln(2)/0/08
t = ~ 8.664 periods
The periods have to match the interest period, ie, 8 years for 8% per year, 8 hours for 8% per hour, etc.
The amount, the $15,000, is not relevant to the problem.